A professor has a 95% passing rate for all of his past classes. He surveys his...

  1. A professor has a 95% passing rate for all of his past classes. He surveys his current 93 students, and 13 of them are worried about failing.

  1. For his class of 93 students, determine the mean and standard deviation of students passing the class.


ive had this question here all day but not help.
Answer #1

Solution :

Given that,

p = 0.95

q = 1 - p = 1 - 0.95 = 0.05

n = 93

Using binomial distribution,

Mean = = n * p = 93 * 0.95 = 88.35

Standard deviation = = n * p * q = 93 * 0.95 * 0.05 = 2.10